OpenStudy (anonymous):

solve for x: 4^(3x+5)=16

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

how do you do it though??

OpenStudy (anonymous):

it's very easy..........

OpenStudy (anonymous):

can u teach me?

OpenStudy (anonymous):

yes, off course , free of cost........

OpenStudy (anonymous):

that's why i m here ......

OpenStudy (anonymous):

lol dont be annoying

OpenStudy (anonymous):

i m not joking..........

OpenStudy (anonymous):

moving on..

OpenStudy (anonymous):

you can just set the equation in the exponent = to 2... becuase u know that 4^2 is 16, so 3x+5=2

OpenStudy (bahrom7893):

Step by step: 4^(3x+5)=16 Rewrite 16 as 4^2: 4^(3x+5) = 4^2 Take Ln of both sides: Ln[4^(3x+5)] = Ln[4^2] Remember the property: LnA^b = b*LnA, so: 3x+5 = 2; 3x = -3; x = -1

OpenStudy (anonymous):

sorry, for late reply, due tp connection problem..